a couple has 3 children. find the probability that the couple has at least one of each gender
Answer by stanbon(75887)

a couple has 3 children. find the probability that the couple has at least one of each gender
-----------
P(at least one boy and one girl) = 1 - [P(no boys)+P(no girls)]
-----
= 1 - [(1/8) + (1/8)]
= 6/8 = 3/4
===================
Cheers,
Stan H.